### 10 Mar 2026 v1

# Omics Data Discovery Agents

Abstract
The biomedical literature contains a vast collection of omics studies, yet most published data remain functionally inaccessible for computational reuse. When raw data are deposited in public repositories, essential information for reproducing reported results is dispersed across main text, supplementary files, and code repositories. In rarer instances where intermediate data is made available (e.g. protein abundance files), its location is irregular. In this article, we present an agentic framework that fetches omics-related articles and transforms the unstructured information into searchable research objects. Our system employs large language model (LLM) agents with access to tools for fetching omics studies, extracting article metadata, identifying and downloading published data, executing containerized quantification pipelines, and running analyses to address novel question. We demonstrate automated metadata extraction from PubMed Central articles, achieving 80% precision for dataset identification from standard data repositories. Using model context protocol (MCP) servers to expose containerized analysis tools, our set of agents were able to identify a set of relevant articles, download the associated datasets, and re-quantify the proteomics data. The results had a 63% overlap in differentially expressed proteins when matching reported preprocessing methods. Furthermore, we show that agents can identify semantically similar studies, determine data compatibility, and perform cross-study comparisons, revealing consistent protein regulation patterns in liver fibrosis. This work establishes a foundation for converting the static biomedical literature into an executable, queryable resource that enables automated data reuse at scale.
Keywords: large language models, omics data, proteomics, transcriptomics, model context protocol, automated curation, data reanalysis, meta-analysis
1 Introduction
The biomedical literature has growing collection of omics studies, yet the data underlying these publications is impractical for computational reuse. Although raw data are increasingly deposited in public repositories that specialize in distributing them, including PRIDE Perez-Riverol and others (2025), MassIVE Wang and others (2018), GEO Edgar (2002); Barrett and others (2012), and the Sequence Read Archive Katz and others (2022). These resources are a critical component for ensuring that the results reported in scientific articles remain verifiable. However, raw data require domain knowledge to ensure that the quantification tools are set up correctly and that acceptable values are used for their parameters. Despite being the most immediately-useful form of the data, the quantified data are often not placed in public repositories and the barriers to reprocess the data significantly discourages data reuse.
When processed quantitative data are provided, they may appear in different locations and with different levels of completeness. Without a standard way of distributing quantitative data, processing parameters and analytical decisions are similarly scattered across the main article text, supplemental materials, and code repositories. Some journals allow authors to make their data available “upon request,” but responses to such requests are frequently incomplete or reveal that the materials are no longer available due to personnel turnover, storage limitations, or similar constraints. A previous study found that data availability declines with article age, with retrieval success rates dropping by 17% each year after publication Vines and others (2014).
The cumulative effect is that most published omics studies cannot be computationally reused without substantial manual effort. In practice, this limits reproducibility, cross-study synthesis, downstream discovery, and interdisciplinary collaboration. The scale of the problem precludes manual solutions: in November 2025 alone, more than 1,500 articles matching the query “proteomics OR mass spectrometry” were available from PubMed Central (PMC). A similar volume of articles are available for genomics, transcriptomics, and metabolomics.
Current large-scale automated curation efforts primarily focus on linking existing author-provided metadata, which is often limited to a small set of keywords. Commercial tools that use AI assistants also exist, but are limited to summarizing literature or generating data-analysis code, rather than identifying datasets and making them accessible via natural language. For instance, Perplexity Perplexity AI (2024) offers literature research but is prone to both direct and second-hand hallucination GPTZero (2024). Platforms such as ScienceMachine ScienceMachine (2024) or Kosmos Mitchener et al. (2025) offer custom analysis of data on an experiment-by-experiment basis but do not leverage the large corpus of public data. At present, there is no scalable system capable of autonomously identifying relevant omics studies, recovering their underlying data and metadata, executing appropriate analyses on demand, and returning auditable, study-level results. Recent advances in large language models (LLMs) and their application to knowledge graph construction Peng et al. (2023) suggest new possibilities for automated literature processing. Emerging work has demonstrated the potential of AI agents for biomedical data analysis Huang and others (2025), though these approaches have not yet been combined with comprehensive literature mining to utilize value of existing biomedical omics publications.
Here, we present an agentic framework that addresses these limitations by automating the identification, extraction, and linking of omics research products from full-text articles and supplemental material. Unlike static metadata-harvesting approaches used by existing repositories, we extract information from the unstructured publication text and link the article to its datasets, sample annotations, processing methods, and code repositories. The structured representation enables searches beyond keyword matching, including semantic queries that better capture article content. Our key innovation is the deployment of AI agents that can locate relevant studies ingested, identify quantitative datasets, fetch and process raw data in a manner similar to publication, and integrate the results of multiple studies. The extracted information allow agents to reason over the full publication record, including the main text, supplementary files, repository metadata, and available code. When processed data is available, we demonstrate that agents are able to identify and locate it whether it is in supplemental materials, standard repositories, or project-specific sites. When only raw data are available, agents infer the processing context from available documentation and invoke standard quantification tools with parameters derived from the publication.
We demonstrate the feasibility of this approach through three core capabilities: (1) automated metadata extraction from articles available via PMC and insertion into a relational database; (2) agent-guided raw data quantification using containerized tools exposed through model context protocol (MCP) servers; and (3) cross-study reasoning that identifies compatible datasets and synthesizes findings across multiple publications. Our results establish that LLM agents can effectively transform unstructured omics publications into reusable, computable research objects. By providing access to containerized tools and published data, we demonstrate that agents can find reproducible results across studies, even when those findings are not reported in the article. Here, we lay the groundwork for executable, agent-driven analysis of the biomedical literature at scale.
2 Methods
2.1 System Architecture Overview
Our agentic curation and reanalysis system consists of three components, shown in Figure 1: (1) an article ingestion and metadata extraction pipeline, (2) a data retrieval and quantification system, (3) a cross-study reasoning system that identifies compatible datasets and synthesizes findings across multiple publications.
Figure 1: Schematic of agent system. Articles and supplemental materials are obtained by the the Article Parsing agent and metadata is extracted into a database. Articles
2.2 Article Ingestion and Metadata Extraction
2.2.1 Data Sources
We retrieved full-text articles from the PubMed Central Open Access (PMC OA) subset using the NCBI E-utilities API. We identified articles returned from the query, "proteomics" for November 2025 - January 2026 in addition to "liver fibrosis proteomics" for 2025, for a total of 4210 articles, with the bulk of the articles coming from the broader "proteomics" search.
2.2.2 LLM-Based Information Extraction
Metadata extraction is done in two steps: gathering existing structured metadata such as title, journal, keywords, etc. that are readily available via PMC XML fields, and gathering unstructured metadata from the full text and supplemental. The full article text was processed using a separate LLM (gpt-5) to extract additional keywords and identify the locations of raw data, analysis code, and quantitative data. The purpose of the separate LLM is to limit the possible effects of prompt injections; the external LLM is instructed to return a JSON-formatted response, and the content is processed by functions that do not expose the contents to the agent. The separation can be achieved via programmatic access to models, which is currently readily available from the major providers (e.g. Anthropic’s Client SDKs, Azure’s AI Foundry) or executed locally via Ollama 15. Agents were instructed to provide evidence citations for all extracted values and to explicitly abstain when information could not be reliably determined. Text embeddings were generated using OpenAI’s text-embedding-3-small model and stored in a database for use with semantic similarity search.
2.3 Evaluation of Metadata Extraction
We manually curated a benchmark set of 39 proteomics publications, annotating all identifiable datasets including raw data repositories, processed data locations, and supplementary materials. Agent extraction results of these articles were compared to manual annotation to compute precision and recall metrics. Results were categorized as correct, incorrect, missed, or ambiguous (where the result is "technically" correct but not semantically so).
2.4 Agent-Guided Raw Data Reanalysis
2.4.1 Model Context Protocol (MCP) Server Design
Analytical tools were made available to agents via function calls provided by MCP servers. This design prevents agents from developing custom tools for each article to ensure reproducibility through containerized execution. MCP servers expose the following method categories:
Data retrieval methods: Transfer data from public data repositories (PRIDE, MassIVE, GEO), inspect file structure and formats.
Quantification pipeline selection and configuration: Using descriptions from the article and available metadata, identify appropriate quantification pipelines (e.g. DIA-NN, MaxQuant) and ensure that assumptions made by pipelines are consistent with data acquisition parameters. Once identified, a configuration file for use in later quantification steps is generated.
Execution methods: Launch containerized analysis pipelines (Apptainer) using information from the previously-generated configuration files. Estimate resource usage and enforce limits.
2.4.2 Containerized Analysis Tools
For proteomics data, the MCP server exposes two containerized quantification tools for the main data acquisition types: DIA-NN Demichev et al. (2020) for data-independent acquisition (DIA) and MaxQuant Tyanova et al. (2016) for data-dependent acquisition (DDA). These workflows produce protein-level abundance matrices mapped to standardized UniProt The UniProt Consortium (2025) gene identifiers.
2.4.3 Agent Workflow for Quantification
When an article is identified for quantification, database tools first verify whether raw data (e.g. mass spectrometry) has been identified. If so, an agent operates the following workflow:
  1. 1.
    Parameter extraction: Parse the article text, supplementary materials, and code repositories to identify analysis parameters including enzyme specificity, missed cleavages, mass tolerances, variable modifications, and software versions.
  2. 2.
    Data retrieval: Use the data retrieval MCP tools to download raw data files from identified repositories.
  3. 3.
    Tool selection: Based on the data acquisition type and article description, select the appropriate containerized analysis tool.
  4. 4.
    Configuration: Generate a configuration file specifying all identified parameters. Parameters not explicitly stated in the article are flagged as defaults.
  5. 5.
    Execution: Invoke the containerized pipeline and monitor for completion or errors.
  6. 6.
    Validation: Compare output dimensions and summary statistics against article-reported values when available.
2.4.4 Evaluation of Quantification
We evaluated reanalysis performance using articles where both raw and processed data were publicly available. Primary metrics included execution success rate, correlation between agent-generated and published protein abundances, and overlap in differentially expressed proteins when statistical analysis was performed.
Figure 2: Comparison of protein identifications reported by Chen et al. and that obtained by ODDA using the article content to inform parameter selection. Filtering: Both datasets were filtered to remove reverse hits, contaminants, and proteins identified only by site. Proteins were matched by UniProt Majority Protein ID and by gene name. The number of proteins with valid intensity values counted for each of the 6 samples (CCl4-1/2/3, Oil-1/2/3). For each sample, proteins present in both datasets were identified and LFQ intensity values were extracted for shared proteins. Intensities were log10-transformed for correlation analysis, then Pearson correlation was calculated per-sample and overall (pooled across all samples).
2.5 Cross-Study Reasoning
2.5.1 Study Compatibility Assessment
For cross-study analyses, an agent was instructed to obtain 3 articles examining similar problems, as determined by the cosine similarity between the text embeddings of their abstracts. When available, the agent was instructed to use already-quantified data. Otherwise, it was to obtain the public raw data and quantify it as described in the corresponding article.
3 Results
3.1 Extracted articles
Of the 4210 articles, 2442 had full-text articles and 2379 had supplemental material available via PMC OA. 1264 (51.8% of full-text articles) articles had references to published raw data, 710 (29.1%) had references to quantitative data, and 464 (19.0%) were found to have both.
3.2 Automated Metadata Extraction from PMC Articles
To evaluate performance, we manually curated a benchmark set of 39 proteomics publications, annotating all identifiable datasets including raw data repositories, processed data, and supplementary materials.
The agent system identified 91 datasets across the 39 articles, with performance metrics summarized in Table 1. When excluding ambiguous cases, the system achieved a precision of 0.91 and recall of 0.89. When classifying ambiguous results as incorrect, precision decreased to 0.53 while recall remained at 0.89.
Table 1: Performance metrics for automated dataset identification from 39 proteomics articles. Results are shown with ambiguous cases omitted and classified as incorrect.
Metric Omit Ambiguous Ambiguous is Incorrect
Correct 91 91
Incorrect 8 81
Missed 11 11
Precision 0.91 0.53
Recall 0.89 0.89
Identification of standard repository links (PRIDE, MassIVE, GEO) was notably more successful, with 49 of 61 correctly identified (approximately 80% precision). Errors included non-existent dataset identifiers (8 cases) and datasets present in the text that the agent did not identify (11 cases). The 73 ambiguous cases typically involved supplementary data files where it was unclear whether the files should be classified as "processed data."
3.3 Agent-Guided Raw Data Quantification
We next evaluated the system’s ability to automatically reprocess raw proteomics data using parameters extracted from article text. We selected the article by Taneera et al. Taneera and others (2026), which describes a DIA proteomics experiment examining the effect of PPP1R1A knockdown on INS-1 cell proteomics.
3.3.1 Parameter Extraction
The agent successfully extracted DIA-NN processing parameters from the article text. Key parameters identified included:
  • Enzyme specificity: Trypsin (mentioned in article)
  • Data acquisition mode: DIA
  • Software: DIA-NN
  • Experimental conditions: Control and PPP knockdown (identified from raw data filenames)
Notably, the agent differentiated between the two experimental conditions (control vs. knockdown) based on filename patterns in the raw data without explicit instruction, demonstrating autonomous interpretation of experimental design when the correspondence between experimental conditions and files is not explicitly stated.
The article reported 9,006 quantified proteins. Initially, only DIA-NN v2.3.1 was made available, and the agent flagged the difference with the article’s reported version (v1.8.1) before proceeding with v2.3.1. After making multiple versions of DIA-NN available, the agent still initially selected the most recent version (v2.3.1), but selected the correct version (v1.8.1) when explicitly prompted to match the analysis specified in the article. The agent identified 9,530 and 7,729 proteins for v1.8.1 and v2.3.1, respectively, which is in coarse agreement with the published results.
3.3.2 Differential Expression Analysis
To further validate the reanalysis, we instructed the agent to investigate differentially expressed proteins (DEPs) between control and PPP1R1A knockdown conditions. The agent correctly identified that the original study used Benjamini-Hochberg correction for multiple testing, extracting this information from the article’s statistical methods section. The initial attempt had the overlap of DEPs at 37%; repeating the process and adding the instruction to reproduce the article’s preprocessing steps increased the overlap to 63%. The substantial increase in DEP overlap indicates that it is important to specify that analyses should reproduce those in an article; otherwise, agents are likely to use the latest available tool. Differences in software versions are known to affect downstream results, so the discrepancy is expected. We further note that Taneera et al. used existing toolkits in R for imputation and differential expression, while the agent used custom Python tools to emulate the behavior. For the purposes of this paper, the 63% overlap was deemed sufficient.
3.4 Cross-Study Compatibility Assessment
We evaluated the system’s ability to identify compatible studies for data integration. When presented with two studies examining proteomics of liver pathology (one one liver cancer, the other on fatty liver disease), the agent correctly determined that the datasets were not directly comparable and declined to perform integration.
When tasked to find 3 similar articles using the minimum pairwise distance measured by the cosine similarity of the abstract embeddings, the agent identified three studies: Cheng et al. Cheng and others (2025), Jirouskova et al. Jirouskova and others (2025), and Devos et al. Devos et al. (2026). All three papers investigated proteomics in liver fibrosis. Cheng et al. examined proteomics and phosphoproteomics to show kinase dysregulation in mice. Jirouskova et al. reported on the role of clusterin, an extracellular chaperone protein. Lastly, Devos et al. reported on differentially expressed proteins in both heart and liver fibrosis in human patients, identifying 18 DEPs shared across the organs. We visualize their proximity using a UMAP (Figure 3).
Figure 3: Distribution of articles in a UMAP. Text embeddings were computed using article abstracts and used to compute the UMAP. The three articles that were identified as semantically similar are highlighted in red.
The agent obtained protein quantities for the raw data shared by Cheng et al. by using MaxQuant as described in the article; differential expression was obtained by comparing across the case/control files. Case vs. control was automatically determined by considering the experimental procedure in the article and the filenames (_CCL4_ vs. _OIL_). The protein quantities for Jirouskova et al. were supplied in the public repository, and those were used for comparison. The 18 DEPs reported by Devos et al. are not listed in the main article, but are found in a nested archive within the supplemental material that was automatically downloaded during article ingestion. When tasked with determining whether the articles demonstrated agreement, the agent obtained the list of 18 proteins from Devos et al. and compared these with the differential expression reported in the data from Cheng et al. and Jirouskova et al., and found that 11/18 proteins showed the same upregulation across species. Of those 11, only 6 were detected across all three experiments: CLU, TGFBI, AMBP, MYH10, PRELP, Col14A1, while the other 5 were detected in only one of the mouse-based experiments. Although the direction of regulation agreed across studies, the degree of regulation was considerably different for most proteins. We note that other than the upregulation of CLU reported as one of the main findings in Jirouskova et al., the remaining proteins are not reported in the main text of any of the three articles. To determine that the three studies demonstrate reproducible differential expression, it was necessary to obtain full article, determine processing parameters, quantify one of the datasets, identify proteins of interest, and compare high-level differential expression for the proteins. Literature review confirmed that all consistently-regulated proteins have previously reported roles in liver fibrosis: CLU Seo et al. (2019); Wang et al. (2026), TGFBI Krzistetzko et al. (2023); Wu et al. (2026), AMBP Hakuno et al. (2018), MYH10Liu et al. (2011), PRELPYamauchi et al. (2024), Col14A1Lai et al. (2011).
3.5 System Performance Summary
Across the evaluated tasks, our agentic system demonstrated:
  1. 1.
    High precision for standard repositories: 80% precision for identifying datasets in PRIDE, MassIVE, and GEO.
  2. 2.
    Successful parameter extraction: Automated identification of processing parameters from article text enabled reproduction of published analyses.
  3. 3.
    Reasonable quantification agreement: Protein counts within expected range of published values; 63% DEP overlap when matching preprocessing.
  4. 4.
    Accurate compatibility assessment: Correct identification of comparable studies and appropriate refusal to integrate incompatible datasets.
  5. 5.
    Biologically valid cross-study integration: Cross-species agreement for differentially-expressed proteins in liver fibrosis.
4 Discussion
4.1 Significance for Omics Data Reuse
The availability of publicly accessible datasets provides an opportunity to verify the results of new experiments using previously published data. While a single article is generally narrow in scope, the associated data may contain information relevant to other questions. Finding relevant data is challenging, especially when only raw data is available, and quantification would otherwise require substantial manual intervention. Agentic systems provide a new opportunity to convert unstructured descriptions of scientific articles into searchable datasets by automating many labor-intensive aspects of data processing.
4.2 Technical Innovations
The main contribution of this manuscript is to demonstrate that datasets contain rich, untapped information that cannot be captured by examining only the article’s main text. Extracting a more-complete metadata requires examining supplemental materials and uploaded datasets. By providing tools accessible via MCP servers for transferring, examining, and processing datasets, agentic systems can be used to obtain that information reliably and transparently. Since most tasks do not require novel analyses (e.g., interacting with APIs), the use of MCP servers reduces the risk that agents execute code with silent errors.
4.3 Limitations
The high-level comparison across studies (e.g. differential expression) requires the execution of experiment-specific code to consider the findings of interest; implementing static MCP-based analyses would require adherence to standardized data structures and would only be applicable to a subset of experiments. Extending agent functionality to multi-study comparisons inherently requires a flexible treatment of experimental data, which in turn introduces the possibility of idiosyncratic analyses.
MCP servers are not ultimate safeguards: they simply provide an interface for pre-defined tools. Malicious or insecure implementations of an MCP server still presents a risk. We note that this is not limited to the system proposed in this article but is a known risk with LLMs using MCP servers.
As with any use of LLMs, prompt injection is a possible attack vector. During testing, using obvious attacks such as local files with fake PMCIDs was detected as an attack when data verification failed (e.g., no article was found for the fake PMCID), and similar for using a real PMCID but with metadata that did not match the article contents. However, adding a prompt into a real article was successful and went unnoticed. For example, including a sentence instructing the system to extract specific keywords successfully introduces the requested keywords into the database. It is unlikely that prompt injection occurs through the contents of a published article, as it would be detected during peer review or after publication. However, the agent system examines not only the article text but also supplementary files, which reviewers are less likely to notice. The risk is mitigated by separating concerns: agents with the ability to use the MCP servers and potentially execute code do not need to read potentially-hostile text and instead use tools that load and send the text to a separate LLM without agency, and the response is processed and stored in the database without the agent seeing it. Prompt injection can still be used to inject classification instructions (e.g. keywords, mislabeling files), but the effects are restricted to the article being processed and is largely inconsequential. The greatest risk is in cross-study synthesis, where the agent inspects the article text to determine how to reproduce the analysis, thereby producing specialized code. Prompt injection at that stage could result in malicious code. We note that the secure way to execute potentially-malicious code is not to do so, and we advise a close inspection of the produced analyses for reasons of both security and scientific understanding.
4.4 Summary
Author Contributions
A.H. and J.G.M. conceived the study. A.H. developed the software, conducted experiments, and wrote the manuscript. A.H. and J.G.M. reviewed the manuscript.
Data Availability
All code for the agentic curation system, MCP servers, and evaluation scripts are available on GitHub. The database schema and containerized tool definitions (Apptainer definition files) are released under GPL v3.0 license. Due to license restrictions on the quantification tools, we are unable to provide the containers directly, but we have included instructions for users on how to obtain the relevant files while respecting license terms.
Competing Interests
The authors declare no competing interests.